OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. express your answer in the form y=mx+b (8,9);x+8y=3

OpenStudy (anonymous):

Alright, first thing we need to do is figure out the slope, since we want a parallel line. Do you recall how to find the slope from an equation?

OpenStudy (anonymous):

y1,y2/x1,x2

OpenStudy (anonymous):

That is a good guess, but that's only true if you just have points. I'll go ahead and tell you. If you have an equation (and not just two points), you find the slope by making the equation say y = (the rest of the stuff). So, how do we make x+8y=3 to where the y is by itself?

OpenStudy (anonymous):

y=-1/8x+10

OpenStudy (anonymous):

Very good try. However, we do it like this: x+8y=3 8y = -x + 3 y = (-x + 3)/8 y = -x/8 + 3/8 Seem legit so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Awesome. And you know that -x/8 can be turned into (-1/8)x, right?

OpenStudy (anonymous):

i have already answered this question correctly

OpenStudy (anonymous):

Ah oh well.

OpenStudy (anonymous):

but thank you

OpenStudy (anonymous):

no prob